### v1

@@ Abstract
This letter proposes the experimental validation of an optimisation method for periodic metasurfaces. A previous study showed the design and simulation of a classical anomalous reflecting metasurface and enlightened the parasitic reflection induced by the translational invariances of the structure. The method applies to periodic structures, in the framework of the Floquet analysis. It consists of the exploitation of the Floquet type simulation outputs of a periodic element of the structure to predict the behaviour of the complete structure in terms of radar cross section (RCS). The proposed application focuses on the reduction of the RCS level of the structure in one particular Floquet direction. The fabricated metasurfaces and experimental setup are presented with the associated measurements. The well-agreeing simulation and experimental results validate the proposed optimization procedure.
@@ I.   Introduction
The radar cross section (RCS) is a measure of the reflective ability of a target [1]. For stealth purposes, one may want to reduce the RCS of an object. A classical way to do this is to shape the object or to use absorbing volumetric materials for its conception. To address this issue, another possibility recently emerged : metasurfaces. They have generated a wide variety of applications for wave front shaping [2]. Among them, the classical gradient metasurface leading to anomalous reflection has been particularly studied [3]. It allows reflecting an incoming illumination in a non specular direction, which may prove useful for stealth purposes. Such a metasurface consists of the juxtaposition of sub-wavelength scattering elements (the cells) introducing a linearly varying phase shift at the surface of the object. However, these classical gradient metasurfaces exhibit parasitic reflection directions because of the periodic nature of the phase gradient [4]. The scatterers used to cover a gradient period form a supercell that, when periodically replicated to pave the overall surface, excites Floquet harmonics [5]. Translational invariances and associated symmetries in metasurfaces are now widely studied [6].
Recently, we proposed a method [7] to mitigate parasitic reflections of an anomalous reflecting metasurface with a monodimensionally varying gradient, using Floquet analysis. Simulating a supercell in a Floquet environment allows the computation of the $S_{m,n}$ scattering parameters between the different Floquet modes. Assuming ${n=0}$ corresponds to the fundamental mode impinging on the metasurface, there is a direct correlation between $S_{m,0}$ and the resulting RCS level in the propagation direction of reflected mode $m$, defined by an angle $\theta_{m}$. This means that the outputs of the Floquet simulation (light to simulate) are sufficient to correctly describe the behaviour of the whole structure (requiring a much heavier simulation) in terms of RCS. Going further in this direction, the Floquet simulation $S$ parameters has been incorporated into an analytical model to calculate the total RCS of the structures. This approach has been successfully validated by simulation in [8]. The purpose of the present letter is to provide a complete experimental demonstration and, more generally speaking, to address practical fabrication and measurement issues.
II.   Fabricated metasurfaces
@@
[…]
@@
III.   Measurement facility
The RCS measurements have been made with a specific 3D measurement facility at the CEA-CESTA, presented in [9]. It consists of an antenna attached on an armature, mounted on a rotating axis. A motor drives the rotation of the whole structure giving access to bistatic RCS measurements. This facility is usually exploited in a monostatic configuration, but, for the current study, an transmitting antenna was added, illuminating the metasurface under normal incidence. The situation is depicted in Figure 3 where both antennas are at a distance ${r=4}$ m from the target.
@@ IV.   Results
Figure 4: Bistatic RCS of both initial and optimised metasurfaces, obtained with a full wave simulation of with RCS measurement.
@@
The measured RCS of the initial and optimised surfaces are represented in Figure 4, with the simulated ones. The measured RCS in the ${\theta=-60^{\circ}}$ direction is 9.05 dB for the initial surface and 2.20 dB for the optimised one. This is a 6.85 dB reduction while the simulation shows a 6.80 dB reduction. The agreement between the simulated reduction and the measured reduction is excellent. Globally, the simulation is overestimating RCS values. Actually, the number of assumptions for the synthesis of the surfaces is important (local periodicity, infinite environment, description with only a phase response, phase response computed at only one frequency). This large amount of assumptions makes the optimisation process very efficient. For the synthesis and simulation steps, every computation is made with far field considerations. From a measurement point of view, with an object that is approximately 40 cm long and an antenna at 4 m from the target, the far field hypothesis is not accurate. This would be the first hypothesis to address in order to explain the simulation to measurement disagreement.
@@
Figure 5 shows the evolution of $\Delta$RCS, the optimisation performance. It is defined as $\Delta\text{RCS}(f,\theta)=\text{RCS}_{\textrm{opti.}}(f,\theta)-\text{RCS}_{\textrm{init.}}(f,\theta)$ where $f$ is the working frequency and $\theta$ is the direction of observation (i.e. the bistatic angle). The evolution of $\Delta\text{RCS}(f,\theta=-60^{\circ})$ is in orange thick line in Figure 5. For clarity purposes the trace domain is restricted to the 6 GHz to 10 GHz frequency band, while the measurement was, as said earlier, conducted in the 2 GHz to 18 GHz band. It is observed that the optimisation is performing in a very narrow frequency band. We report a 3 dB reduction in a thin 0.2 GHz band, that is represented by a grey rectangle in Figure 5. A large region around 9 GHz interestingly shows a very good performance over a 1 GHz band. It should be kept in mind that no design rule has been followed toward the synthesis of a wide band device. More generally, metasurfaces devices fail to handle wideband phenomena, or at a cost of sophisticated synthesis procedures (multi scales or multi layer devices [10]) that are not practical for stealth applications. For completeness, $\Delta\text{RCS}(f,\theta=\theta_{-1})$ has been calculated, where $\theta_{-1}$ is the $m=-1$ Floquet mode direction which varies with frequency. This curve is shown as a thin grey line in the same figure. The trace shows a similar behaviour around the central frequency.
@@
V.   Conclusion
This study demonstrates the relevance of the optimization method by proposing a measurement of an optimised surface. The simulation and measurement exhibit a RCS reduction of almost 7 dB, in the direction of the parasitic lobe, with an excellent agreement between them. It has been made possible by the simplicity of the Floquet type simulation and the correct use of its outputs to predict the behaviour of the surface. The method should not be limited to the RCS reduction in a particular direction. One could consider applying this method for any power repartition, in other directions or multiple ones, as long as one stays in the framework of the Floquet analysis.

### v2

@@ Abstract
This letter proposes the design and the measurement of a periodic metasurface that performs anomalous reflection with reduced RCS in a parasitic direction. A previous study proposed a semi-analytical model to predict the behaviour, in terms of RCS, of such a metasurface. This work will therefore focus the discussion on fabrication and measurement issues. The synthesis, design specifications, fabrication method and experimental setup are presented and discussed. The measurement results are extensively discussed as well, enlightening some limitations in the context of metasurface RCS measurements. The proposed metasurface indeed performs the anticipated reduction of the RCS level in the considered Floquet direction. The well-agreeing simulation and experimental results demonstrates the accuracy of the modelling and the efficiency of the optimisation procedure.
@@ I.   Introduction
The radar cross section (RCS) is a measure of the reflective ability of a target [Knott2004]. For stealth purposes, one may want to reduce the RCS of an object. A classical way to do this is to shape the object or to use absorbing volumetric materials for its conception. To address this issue, another possibility recently emerged : metasurfaces. They have generated a wide variety of applications for wave front shaping [Glybovski2016]. Among them, the classical gradient metasurface leading to anomalous reflection has been particularly studied [Yu2011]. It allows reflecting an incoming illumination in a non specular direction, which may prove useful for stealth purposes. Such a metasurface consists of the juxtaposition of sub-wavelength scattering elements (the cells) introducing a linearly varying phase shift at the surface of the object. However, these classical gradient metasurfaces exhibit parasitic reflection directions because of the periodic nature of the phase gradient [DiazRubio2017]. The scatterers used to cover a gradient period form a supercell that, when periodically replicated to pave the overall surface, excite Floquet harmonics [Bhattacharyya2006]. Such translational invariances and associated symmetries in metasurfaces are now widely studied [QuevedoTeruel2020].
Recently, we proposed a method [Elineau2022a] to mitigate parasitic reflections of an anomalous reflecting metasurface with a monodimensionally varying gradient, using Floquet analysis. Simulating a supercell in a Floquet environment allows the computation of the $S_{m,n}$ scattering parameters between the different Floquet modes. Assuming ${n=0}$ corresponds to the fundamental mode impinging on the metasurface, there is a direct correlation between $S_{m,0}$ and the resulting RCS level in the propagation direction of reflected mode $m$, defined by an angle $\theta_{m}$. This means that the outputs of the Floquet simulation (light to simulate) may be sufficient to correctly describe the behaviour of the whole structure (requiring a much heavier simulation) in terms of RCS. Going further in this direction, the Floquet simulation $S$ parameters has been incorporated into an analytical model to calculate the total RCS of the structures. This approach has been presented in [Elineau2022b] proving that the outputs of the Floquet type simulation are indeed sufficient to estimate the RCS of the metasurfaces. The purpose of the present letter is to conclude the demonstration by providing the design and measurement of such a metasurface and, more generally speaking, to address the associated practical fabrication and measurement issues.
II.   Metasurfaces synthesis
@@
An optimised surface is proposed by following the optimisation procedure of [Elineau2022b]. The optimisation is performed at a unique frequency ${f=8}$ GHz and consists of the maximisation of the difference between the $S_{1,0}$ and $S_{-1,0}$ parameters, using cell dimensions as degrees of freedom. This process gives an optimised supercell whose dimensions are also listed in figure 1. The use of this optimised supercell results in a 6.80 dB reduction of the RCS level in the ${\theta_{-1}=-60^{\circ}}$ direction, as shown in figure 4. Dashed line corresponds to the initial surface and plain line corresponds to the optimised one.
III.   Metasurfaces fabrication
Both surfaces, initial and optimised are fabricated with printed copper patches on a Neltec NY9217 copper backed substrate. The relative permittivity of such a substrate is low ($\epsilon_{r}=2.17$) which provides smooth varying phase variations of the cells with their geometrical parameter as well as with $\theta$ variations. This helps under these two aspects, at least. In one hand, finding converging optimisations is made easier and, in the other hand, measurements are made less sensitive to $\theta$ uncertainties. A ${27\times 27}$ cells surface is fabricated and gives an array of approximately ${10\lambda\times 10\lambda}$ (or approximately ${40\times 40}$ cm2). A close view of each surface is found in Figure 2 where a supercell is enclosed by a dashed line.
@@ III.   Metasurfaces fabrication
IV.   Measurement facility
The RCS measurements have been made with a specific 3D measurement facility at the CEA-CESTA, presented in [Massaloux2014]. It consists of an antenna attached on an armature, mounted on a rotating axis. A motor drives the rotation of the whole structure giving access to bistatic RCS measurements. This facility is usually exploited in a monostatic configuration, but, for the current study, a transmitting antenna was added, illuminating the metasurface under normal incidence. The situation is depicted in Figure 3 where both antennas are at a distance ${r=4}$ m from the target.
@@ V.   Results
Figure 4: Bistatic RCS of both initial and optimised metasurfaces, obtained with a full wave simulation or with RCS measurement.
@@
The measured RCS of the initial and optimised surfaces are represented in Figure 4, with the simulated ones. The measured RCS in the ${\theta=-60^{\circ}}$ direction is 9.05 dB for the initial surface and 2.20 dB for the optimised one. This is a 6.85 dB reduction while the simulation shows a 6.80 dB reduction. The agreement between the simulated reduction and the measured reduction is excellent. Globally, the simulation is overestimating RCS values. Actually, the number of assumptions for the synthesis of the surfaces is important (local periodicity, infinite environment, description with only a phase response, phase response computed at only one frequency). This large amount of assumptions lighter the optimisation process to make it very efficient.
For the synthesis and simulation steps, every computation is made with far field considerations. From a measurement point of view, with an object that is approximately 40 cm long and an antenna at 4 m from the target, the far field hypothesis is not accurate. It is possible to estimate some consequences of the far field hypothesis violation during measurements. The phase difference between the centre and an edge of the surface can easily be estimated. In the configuration of the experimental setup presented in figure 3 with a wave number $k$ associated with the frequency ${f=8}$ GHz, we found an approximately $\pi$ phase difference. In contrast, the phase response variations with $L$ of a cell as observed in [Elineau2022b] is of 280.mm-1 at its maximum. For a fabrication tolerance of 50 µm this gives, in the worst case, an uncertainty of 14 which is way under the $\pi$ calculated phase difference.
@@
Figure 5 shows the evolution of $\Delta$RCS, the optimisation performance. It is defined as $\Delta\text{RCS}(f,\theta)=\text{RCS}_{\textrm{opti.}}(f,\theta)-\text{RCS}_{\textrm{init.}}(f,\theta)$ where $f$ is the working frequency and $\theta$ is the direction of observation (the bistatic angle). The evolution of $\Delta\text{RCS}(f,\theta=-60^{\circ})$ is in orange thick line in Figure 5. For clarity purposes the trace domain is restricted to the 6 GHz to 10 GHz frequency band, while the measurement was, as said earlier, conducted in the 2 GHz to 18 GHz band. It is observed that the optimisation is performing in a very narrow frequency band. We report a 3 dB reduction in a thin 0.2 GHz band, that is represented by a grey rectangle in Figure 5. A large region around 9 GHz interestingly shows a very good performance over a 1 GHz band. It should be kept in mind that no design rule has been followed toward the synthesis of a wide band device. More generally, metasurfaces devices fail to handle wideband phenomena, or at a cost of sophisticated synthesis procedures (multi scales or multi layer devices [Samadi2021]) that are not practical for stealth applications. For completeness, $\Delta\text{RCS}(f,\theta=\theta_{-1})$ has been calculated, where $\theta_{-1}$ is the $m=-1$ Floquet mode direction which varies with frequency. This curve is shown as a thin grey line in the same figure. The trace shows a similar behaviour around the central frequency.
@@
VI.   Conclusion
This study demonstrates the efficiency of the RCS reduction metasurface optimisation method by proposing a measurement of such a surface. The simulation and measurement exhibit a RCS reduction of almost 7 dB, in the direction of the parasitic lobe, with an excellent agreement between them. It has been made possible by the simplicity of the Floquet type simulation and the correct use of the outputs of these simulations to predict the behaviour of the surface. The method should not be limited to the RCS reduction in a particular direction. One could consider applying this method for any Floquet mode distribution, addressing one or multiple modes. The literature provides various recent analysis of multichannel Floquet metasurfaces (in [Vuyyuru2024] for example), but none of them approaches the problem from a RCS point of view. We think that this end to end study built the bridge between Floquet analysis and RCS estimation. Multichannel RCS reduction should not show more limitations than the ones already coming from classical multichannel Floquet metasurfaces.